OpenStudy (anonymous):

Find an equation of the line passing through the two given points (1,-12) and (-2,-3)

OpenStudy (anonymous):

y = ax + b is the form of this equation. a is the slope, found by (y2-y1)/(x2-x1) b is the y-intercept. found when x = 0.

OpenStudy (anonymous):

Slope is (-12-(-3))/1-(-2) = -9/3 = -3 Line with this slope using point (-2,-3) is y = -3(x-(-2)) -3 y = -3x-9 = -3(x+3)

OpenStudy (anonymous):

Where x1, y1 is a known point and m is the gradient. \[y - y_{1} = m(x - x_{1})\] \[y - (-12) = -3(x - 1)\] \[y = -3x - 9\]